OpenStudy (anonymous):

solve the equation x+4 / 2 - x-4/4 = 4

OpenStudy (anonymous):

4((x+4)/2 - (x-4)/4) = 16 2(x+4) - (x-4) = 16 2x+8 - x +4 = 16 x + 12 = 16 x = 4 check it -> (4+4)/2 = 4, (4-4)/4 = 0, 4-0 = 4!